# TempR1: Improving Temporal Understanding of MLLMs via Temporal-Aware Multi-Task Reinforcement Learning

### v2

@@ 1 Introduction
Understanding temporal dynamics and semantics in videos is fundamental for comprehensive video comprehension. Temporal video understanding encompasses a variety of tasks that require models to reason about when and how events occur and evolve over time., such as Temporal Grounding (TG) [charadesta, anetgrounding], Dense Temporal Grounding (DTG) [anetgrounding, densegrounding], Temporal Action Localization (TAL) [thumos, activitynet], Video Highlight Detection (VHD) [qvhighlights], and Grounded Video Question Answering (GVQA) [nextgqa]. This capability is essential for real-world applications including intelligent video retrieval, human–computer interaction, and long-form video analysis.
@@
The emergence of MLLMs [gemini, qwen25vl, gpt4, llava] has shifted research toward end-to-end multimodal reasoning, enabling unified video–text processing within a single architecture. Beyond architectural unification, MLLMs naturally support knowledge transfer across diverse tasks and datasets, opening new opportunities for multi-task and cross-domain temporal understanding without the need for handcrafted task-specific designs. Current MLLM-based approaches for temporal understanding mainly fall into two paradigms: supervised fine-tuning (SFT) and reinforcement learning (RL).
SFT-based methods [timesuite, trace, vtimellm, timechat] enhance temporal comprehension through large-scale instruction tuning, but they often suffer from overfitting on limited temporal datasets and a loss of general reasoning ability due to rigid supervision [timer1, videochat-r1]. In contrast, RL-based approaches [videochat-r1, timer1, tar-tvg, museg] directly optimize task-specific objectives and reward temporally and semantically plausible predictions, leading to better data efficiency and generalization. However, most existing RL methods remain constrained to a narrow scope of temporal understanding, typically temporal grounding, which limits their generalization to broader reasoning scenarios such as dense grounding, action localization, and time-sensitive question answering. Moreover, their lack of exposure to diverse temporal structures prevents them from capturing the hierarchical and compositional nature of temporal dependencies.
@@ 2 Related Work
Video Temporal Understanding. Video temporal understanding [charadesta, anetgrounding, activitynet, densegrounding, nextgqa, qvhighlights] aims to analyze the temporal structure and semantic dynamics of videos. Early approaches [umt, qd-detr, univtg, eatr, momentdiff, tr-detr, cg-detr, uvcom, r2tuning, flashvtg, prvg, hscnet] primarily rely on VLP models [clip, videoclip, internvideo, internvideo2, unmasked] to extract vision and text representations, followed by cross-modal attention and task-specific heads for objective prediction. While effective, these task-specific architectures require distinct expert models for each dataset, limiting scalability and cross-task generalization.
Recent advances in MLLMs [gemini, qwen25vl, gpt4, llava, videochat, mvbench, videochat-online, llava-onevision, video-llava, video-llama] have demonstrated strong open-vocabulary understanding and multimodal reasoning capabilities. However, due to limitations in pretraining objectives and data, they still struggle to capture fine-grained temporal dependencies and align linguistic semantics with temporal visual cues. Supervised fine-tuning (SFT) has been explored to enhance temporal comprehension [chatvtg, wang2024efficient, groundinggpt, trace, moviechat, timesuite, vtimellm, timechat, hawkeye, etbench]. Yet, its rigid token-level supervision tends to cause overfitting to instruction data.
In contrast, RL-based fine-tuning directly optimizes task-specific objectives and rewards all reasonable predictions, improving data efficiency and generalization. Recent RL approaches [videochat-r1, timer1, museg, tar-tvg, videochat-r15] achieve strong performance on individual temporal understanding tasks, but mostly remain limited to temporal grounding and lack a unified framework for broader temporal tasks. In this work, we extend RL-based temporal understanding to multiple tasks and propose a scalable reward formulation that adapts to diverse temporal correspondence patterns. By jointly training across these tasks, our proposed TempR1 achieves consistent improvements and establishes a generalizable paradigm for enhancing MLLMs’ temporal understanding.
MLLM Reinforcement Fine-tuning. Following the success of RL in enhancing the reasoning abilities of LLMs through verifiable reward mechanisms [openaio1, deepseekr1], recent works have extended RL optimization to MLLMs. These methods [shen2025vlm, liu2025visual, chen2025visrl, xia2025visionary, huang2025vision, wang2025videorft, feng2025video, bai2025univg, tang2025tspo, zhang2025tinyllava, ouyang2025spacer, chen2025scaling, sc-captioner] design task-specific, rule-based rewards to strengthen multimodal reasoning, leading to improvements in perception, visual grounding, and compositional understanding. However, most of them focus on static images or short video clips, with limited exploration of long-form temporal reasoning tasks beyond temporal grounding.
@@ 3.1 Revisiting Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) [deepseekr1] extends Proximal Policy Optimization (PPO) by replacing critic-based value estimation with group-wise relative comparison, eliminating the need for a separate critic network and greatly reducing training overhead. Specifically, given an input prompt $p$, the policy model $\pi_{\theta}$ samples a group of $G$ candidate responses $o=\{o_{1},\dots,o_{G}\}$. Each response is evaluated by a reward function $r(\cdot)$, yielding the corresponding scores $\{r(o_{1}),\dots,r(o_{G})\}$. Instead of learning an absolute value function, GRPO normalizes these scores within the group to obtain relative advantages that reflect the comparative quality of each response:
@@ 3.3 Temporal Localization Reward
VLP M-DETR [qvhighlights] - - 55.7 34.2 - - - - - - 53.9 34.8
UMT [umt] - - 49.4 26.2 - - - - - - 60.3 44.3
QD-DETR [qd-detr] - - 57.3 32.6 - - - - - - 62.7 46.7
UniVTG [univtg] 52.2 72.6 60.2 38.6 - 56.1 43.4 24.3 - - 59.7 -
SSRN [ssrn] - - 65.6 42.7 - - 54.5 33.2 - - - -
EaTR [eatr] - - 68.5 44.9 - - 58.2 37.6 - - 61.4 45.8
CG-DETR [cg-detr] 50.1 70.4 58.4 36.3 - - - - - - 67.4 52.1
FlashVTG [flashvtg] - - 70.3 49.9 - - - - - - 73.1 57.3
SFT ChatVTG [chatvtg] 34.9 52.7 33.0 15.9 27.2 40.7 22.5 9.4 - - - -
TimeChat [timechat] - - 32.2 13.4 - - - - - - - -
HawkEye [hawkeye] 33.7 50.6 31.4 14.5 32.7 49.1 29.3 10.7 - - - -
HawkEye* [hawkeye] 49.3 72.5 58.3 28.8 39.1 55.9 34.7 17.9 - - - -
VTimeLLM [vtimellm] 31.2 51.0 27.5 11.4 30.4 44.0 27.8 14.3 - - - -
TimeSuite [timesuite] - 69.9 48.7 24.0 - - - - - - - -
TimeSuite* [timesuite] - 79.4 67.1 43.0 - - - - - - - -
TRACE [trace] - - 40.3 19.4 39.0 - 37.7 24.0 - - - -
Qwen-2.5-VL-7B [qwen25vl] 29.6 46.1 25.5 11.5 17.2 22.5 13.0 6.3 12.9 11.8 6.3 3.0
RL Time-R1* [timer1] 58.8 82.8 72.2 50.1 - 73.3 55.6 34.0 44.7 63.2 47.1 28.0
TAR-TVG* [tar-tvg] 61.1 83.6 71.4 50.2 - - - - 65.9 85.6 76.1 58.5
TAR-TVG [tar-tvg] - - - - 41.1 61.5 39.8 19.8 - - - -
VideoChat-R1 [videochat-r1] 60.8 - 71.7 50.2 36.6 - 33.4 17.7 50.1 70.5 53.7 34.1
VideoChat-R1.5 [videochat-r15] 60.6 82.8 71.6 48.3 20.2 28.3 15.6 7.2 18.0 24.5 11.7 4.3
MUSEG [museg] 59.7 - - - 26.5 38.3 22.4 10.5 33.5 49.4 30.8 15.5
@@ 4.1 Training Data
We build a large-scale multi-task corpus for temporal video understanding, containing about 60K video samples collected from multiple datasets. Specifically, TG data come from Charades-STA [charadesta], DiDeMo [didemo], and TimeRFT [timer1]; DTG from ActivityNet-Caption [anetgrounding]; VHD from QVHighlights [qvhighlights]; GVQA from NExTGQA [nextgqa]; and TAL from ActivityNet-v1.3 [activitynet] and HACS [hacs].
@@
VLP PRVG [prvg] 55.6 78.3 61.2 37.8 - - - - - -
HSCNet [hscnet] 59.7 81.9 66.6 44.0 - - - - - -
MLLM Qwen2.5-VL-7B [qwen25vl] 14.1 18.3 10.3 5.0 15.3 20.6 11.1 5.2 60.8 35.9
Time-R1 [timer1] 33.0 48.5 29.7 13.9 28.3 39.2 18.6 8.2 62.1 52.8
VideoChat-R1 [videochat-r1] 34.3 50.7 31.3 15.3 36.1 - - - 70.6 55.6
VideoChat-R1.5 [videochat-r15] 14.6 20.0 10.9 5.2 20.5 28.2 12.7 4.6 62.4 26.8
MUSEG [museg] 32.6 48.5 29.5 13.5 25.2 34.7 16.4 7.4 63.5 58.0
@@ 4.2 Benchmarks and Evaluation Metrics
We evaluate TempR1 on five temporal understanding tasks using standard public benchmarks. For TG, evaluations are conducted on Charades-STA [charadesta] and ActivityNet-Caption [anetgrounding] using mean IoU (mIoU) and Recall at various IoU thresholds. DTG is evaluated on ActivityNet-Caption [anetgrounding] with the same metrics. VHD is assessed on QVHighlights [qvhighlights] following TAR-TVG [tar-tvg], where multiple highlight segments are aggregated for IoU computation, and both mIoU and Recall are reported. GVQA is evaluated on NExT-GQA [nextgqa], reporting answer accuracy and the mean IoU of supporting evidence. TAL follows MUSEG [museg] and uses the mean F1 score (mF1) under four IoU thresholds (0.1, 0.3, 0.5, 0.7). To further assess general understanding, we also test on VideoMME [videomme], Perception Test [perceptiontest], TempCompass [tempcompass], and MVBench [mvbench].
@@ 4.3 Implementation Details
We adopt Qwen2.5-VL-7B [qwen25vl] as the base model for reinforcement fine-tuning. Following existing works [museg], video frames are sampled at 2 FPS, and when the total number of frames exceeds 448, we uniformly sample 448 frames. The spatial resolution of the sampled frames is dynamically scaled so that the total number of visual tokens does not exceed 3,584. During training, we mix data from all five temporal understanding tasks and randomly sample the data batch at each training step. For samples within a batch, task-specific prompts and reward functions are selected according to their task type. The model is trained for one epoch on the 60K multi-task training corpus. During inference, we apply the same frame sampling strategy and task-specific textual prompts as in training. The model outputs are parsed using regular expression matching to extract temporal intervals and structured results.
@@
DAPO [dapo] 60.0 72.0 48.8 60.3 67.7 43.8 36.9 33.7 15.7 69.4 69.5 76.2 58.3 47.6
GSPO [gspo] 82.5 72.3 50.8 59.9 66.6 43.5 38.4 34.8 17.0 70.7 72.6 80.6 63.4 71.7
GRPO [deepseekmath] 61.4 72.7 51.2 59.8 66.8 43.4 39.2 35.9 17.1 70.1 71.1 79.4 61.4 71.0
@@ 4.4 Comparison with State-of-the-arts
We compare our TempR1 against both expert VLP-based methods and recent MLLM-based methods across five temporal understanding tasks, as reported in Tables 1 and 2. As shown in Table 1, TempR1 achieves state-of-the-art performance on Charades-STA [charadesta] and ActivityNet-Caption [anetgrounding] for TG, as well as on QVHighlights [qvhighlights] for VHD. Through multi-task joint training, TempR1 attains 61.4 mIoU on Charades-STA, surpassing VideoChat-R1 [videochat-r1] and MUSEG [museg] by 0.6 and 1.7 points, respectively. On the more challenging QVHighlights dataset, which involves multi-segment retrieval, TempR1 shows a significantly larger gain, reaching 71.1 mIoU and outperforming the second-best TAR-TVG [tar-tvg] by 5.2 mIoU. Remarkably, this performance is achieved after training for only one epoch on the combined multi-task dataset. With additional dataset-specific fine-tuning, TempR1 further improves to 62.5 mIoU on Charades-STA (surpassing TAR-TVG by 1.4 points), 51.9 mIoU on ActivityNet, and 71.6 mIoU on QVHighlights.
As shown in Table 2, TempR1 also excels on more complex tasks, including DTG, GVQA, and TAL. It achieves 59.8 mIoU on ActivityNet [anetgrounding] for DTG, comparable to the best VLP expert HSCNet [hscnet]. On NExTGQA [nextgqa], TempR1 obtains 70.1% question-answering accuracy and 39.2 mIoU for visual evidence localization, improving temporal localization by 3.1 points over VideoChat-R1 [videochat-r1] while maintaining similar answer accuracy. For TAL on ActivityNet-v1.3 [activitynet], TempR1 achieves 71.0 mF1, outperforming MUSEG [museg] by a substantial 13.0 points, validating the effectiveness of our prediction-groundtruth matching strategy and reward design.
We further evaluate TempR1 on general video understanding benchmarks, including VideoMME [videomme], Perception Test [perceptiontest], TempCompass [tempcompass], and MVBench [mvbench], as illustrated in Figure 3. Using the same training data, SFT tends to weaken general reasoning ability compared with the base model, while our reinforcement fine-tuning notably enhances it. This trend aligns with prior observations in VideoChat-R1 [videochat-r1] and Time-R1 [timer1]. Overall, TempR1 delivers the best comprehensive performance across diverse temporal understanding tasks without compromising general video comprehension. These results demonstrate that reinforcement fine-tuning with scalable, task-aligned rewards substantially enhances both temporal comprehension and generalization capabilities of MLLMs in diverse video understanding scenarios.

### v3

# TempR1: Improving Temporal Understanding of MLLMs via Time-Aware Multi-Task Reinforcement Learning

@@ 1 Introduction
Understanding temporal dynamics and semantics in videos is fundamental for comprehensive video comprehension. Temporal video understanding encompasses a variety of tasks that require models to reason about when and how events occur and evolve over time., such as Temporal Grounding (TG) [12, 22], Dense Temporal Grounding (DTG) [22, 5], Temporal Action Localization (TAL) [19, 6], Video Highlight Detection (VHD) [23], and Grounded Video Question Answering (GVQA) [62]. This capability is essential for real-world applications including intelligent video retrieval, human–computer interaction, and long-form video analysis.
@@
The emergence of MLLMs [54, 3, 1, 31] has shifted research toward end-to-end multimodal reasoning, enabling unified video–text processing within a single architecture. Beyond architectural unification, MLLMs naturally support knowledge transfer across diverse tasks and datasets, opening new opportunities for multi-task and cross-domain temporal understanding without the need for handcrafted task-specific designs. Current MLLM-based approaches for temporal understanding mainly fall into two paradigms: supervised fine-tuning (SFT) and reinforcement learning (RL).
SFT-based methods [67, 15, 16, 44] enhance temporal comprehension through large-scale instruction tuning, but they often suffer from overfitting on limited temporal datasets and a loss of general reasoning ability due to rigid supervision [56, 27]. In contrast, RL-based approaches [27, 56, 13, 37] directly optimize task-specific objectives and reward temporally and semantically plausible predictions, leading to better data efficiency and generalization. However, most existing RL methods remain constrained to a narrow scope of temporal understanding, typically temporal grounding, which limits their generalization to broader reasoning scenarios such as dense grounding, action localization, and time-sensitive question answering. Moreover, their lack of exposure to diverse temporal structures prevents them from capturing the hierarchical and compositional nature of temporal dependencies.
@@ 2 Related Work
Video Temporal Understanding. Video temporal understanding [12, 22, 6, 5, 62, 23] aims to analyze the temporal structure and semantic dynamics of videos. Early approaches [33, 39, 30, 21, 26, 49, 38, 63, 32, 7, 47, 50] primarily rely on VLP models [43, 64, 58, 57, 25] to extract vision and text representations, followed by cross-modal attention and task-specific heads for objective prediction. While effective, these task-specific architectures require distinct expert models for each dataset, limiting scalability and cross-task generalization.
Recent advances in MLLMs [54, 3, 1, 31, 52, 51, 18, 24, 29, 68] have demonstrated strong open-vocabulary understanding and multimodal reasoning capabilities. However, due to limitations in pretraining objectives and data, they still struggle to capture fine-grained temporal dependencies and align linguistic semantics with temporal visual cues. Supervised fine-tuning (SFT) has been explored to enhance temporal comprehension [42, 60, 28, 15, 48, 67, 16, 44, 59, 34]. Yet, its rigid token-level supervision tends to cause overfitting to instruction data.
In contrast, RL-based fine-tuning directly optimizes task-specific objectives and rewards all reasonable predictions, improving data efficiency and generalization. Recent RL approaches [27, 56, 37, 13, 65] achieve strong performance on individual temporal understanding tasks, but mostly remain limited to temporal grounding and lack a unified framework for broader temporal tasks. In this work, we extend RL-based temporal understanding to multiple tasks and propose a scalable reward formulation that adapts to diverse temporal correspondence patterns. By jointly training across these tasks, our proposed TempR1 achieves consistent improvements and establishes a generalizable paradigm for enhancing MLLMs’ temporal understanding.
MLLM Reinforcement Fine-tuning. Following the success of RL in enhancing the reasoning abilities of LLMs through verifiable reward mechanisms [20, 14], recent works have extended RL optimization to MLLMs. These methods [46, 36, 9, 61, 17, 55, 10, 4, 53, 70, 40, 8, 69] design task-specific, rule-based rewards to strengthen multimodal reasoning, leading to improvements in perception, visual grounding, and compositional understanding. However, most of them focus on static images or short video clips, with limited exploration of long-form temporal reasoning tasks beyond temporal grounding.
@@ 3.1 Revisiting Group Relative Policy Optimization
Group Relative Policy Optimization (GRPO) [14] extends Proximal Policy Optimization (PPO) by replacing critic-based value estimation with group-wise relative comparison, eliminating the need for a separate critic network and greatly reducing training overhead. Specifically, given an input prompt $p$, the policy model $\pi_{\theta}$ samples a group of $G$ candidate responses $o=\{o_{1},\dots,o_{G}\}$. Each response is evaluated by a reward function $r(\cdot)$, yielding the corresponding scores $\{r(o_{1}),\dots,r(o_{G})\}$. Instead of learning an absolute value function, GRPO normalizes these scores within the group to obtain relative advantages that reflect the comparative quality of each response:
@@ 3.3 Temporal Localization Reward
VLP M-DETR [23] - - 55.7 34.2 - - - - - - 53.9 34.8
UMT [33] - - 49.4 26.2 - - - - - - 60.3 44.3
QD-DETR [39] - - 57.3 32.6 - - - - - - 62.7 46.7
UniVTG [30] 52.2 72.6 60.2 38.6 - 56.1 43.4 24.3 - - 59.7 -
SSRN [73] - - 65.6 42.7 - - 54.5 33.2 - - - -
EaTR [21] - - 68.5 44.9 - - 58.2 37.6 - - 61.4 45.8
CG-DETR [38] 50.1 70.4 58.4 36.3 - - - - - - 67.4 52.1
FlashVTG [7] - - 70.3 49.9 - - - - - - 73.1 57.3
SFT ChatVTG [42] 34.9 52.7 33.0 15.9 27.2 40.7 22.5 9.4 - - - -
TimeChat [44] - - 32.2 13.4 - - - - - - - -
HawkEye [59] 33.7 50.6 31.4 14.5 32.7 49.1 29.3 10.7 - - - -
HawkEye* [59] 49.3 72.5 58.3 28.8 39.1 55.9 34.7 17.9 - - - -
VTimeLLM [16] 31.2 51.0 27.5 11.4 30.4 44.0 27.8 14.3 - - - -
TimeSuite [67] - 69.9 48.7 24.0 - - - - - - - -
TimeSuite* [67] - 79.4 67.1 43.0 - - - - - - - -
TRACE [15] - - 40.3 19.4 39.0 - 37.7 24.0 - - - -
Qwen-2.5-VL-7B [3] 29.6 46.1 25.5 11.5 17.2 22.5 13.0 6.3 12.9 11.8 6.3 3.0
RL Time-R1* [56] 58.8 82.8 72.2 50.1 - 73.3 55.6 34.0 44.7 63.2 47.1 28.0
TAR-TVG* [13] 61.1 83.6 71.4 50.2 - - - - 65.9 85.6 76.1 58.5
TAR-TVG [13] - - - - 41.1 61.5 39.8 19.8 - - - -
VideoChat-R1 [27] 60.8 - 71.7 50.2 36.6 - 33.4 17.7 50.1 70.5 53.7 34.1
VideoChat-R1.5 [65] 60.6 82.8 71.6 48.3 20.2 28.3 15.6 7.2 18.0 24.5 11.7 4.3
MUSEG [37] 59.7 - - - 26.5 38.3 22.4 10.5 33.5 49.4 30.8 15.5
@@ 4.1 Training Data
We build a large-scale multi-task corpus for temporal video understanding, containing about 60K video samples collected from multiple datasets. Specifically, TG data come from Charades-STA [12], DiDeMo [2], and TimeRFT [56]; DTG from ActivityNet-Caption [22]; VHD from QVHighlights [23]; GVQA from NExTGQA [62]; and TAL from ActivityNet-v1.3 [6] and HACS [71].
@@
VLP PRVG [47] 55.6 78.3 61.2 37.8 - - - - - -
HSCNet [50] 59.7 81.9 66.6 44.0 - - - - - -
MLLM Qwen2.5-VL-7B [3] 14.1 18.3 10.3 5.0 15.3 20.6 11.1 5.2 60.8 35.9
Time-R1 [56] 33.0 48.5 29.7 13.9 28.3 39.2 18.6 8.2 62.1 52.8
VideoChat-R1 [27] 34.3 50.7 31.3 15.3 36.1 - - - 70.6 55.6
VideoChat-R1.5 [65] 14.6 20.0 10.9 5.2 20.5 28.2 12.7 4.6 62.4 26.8
MUSEG [37] 32.6 48.5 29.5 13.5 25.2 34.7 16.4 7.4 63.5 58.0
@@ 4.2 Benchmarks and Evaluation Metrics
We evaluate TempR1 on five temporal understanding tasks using standard public benchmarks. For TG, evaluations are conducted on Charades-STA [12] and ActivityNet-Caption [22] using mean IoU (mIoU) and Recall at various IoU thresholds. DTG is evaluated on ActivityNet-Caption [22] with the same metrics. VHD is assessed on QVHighlights [23] following TAR-TVG [13], where multiple highlight segments are aggregated for IoU computation, and both mIoU and Recall are reported. GVQA is evaluated on NExT-GQA [62], reporting answer accuracy and the mean IoU of supporting evidence. TAL follows MUSEG [37] and uses the mean F1 score (mF1) under four IoU thresholds (0.1, 0.3, 0.5, 0.7). To further assess general understanding, we also test on VideoMME [11], Perception Test [41], TempCompass [35], and MVBench [51].
@@ 4.3 Implementation Details
We adopt Qwen2.5-VL-7B [3] as the base model for reinforcement fine-tuning. Following existing works [37], video frames are sampled at 2 FPS, and when the total number of frames exceeds 448, we uniformly sample 448 frames. The spatial resolution of the sampled frames is dynamically scaled so that the total number of visual tokens does not exceed 3,584. During training, we mix data from all five temporal understanding tasks and randomly sample the data batch at each training step. For samples within a batch, task-specific prompts and reward functions are selected according to their task type. The model is trained for one epoch on the 60K multi-task training corpus. During inference, we apply the same frame sampling strategy and task-specific textual prompts as in training. The model outputs are parsed using regular expression matching to extract temporal intervals and structured results.
@@
DAPO [66] 60.0 72.0 48.8 60.3 67.7 43.8 36.9 33.7 15.7 69.4 69.5 76.2 58.3 47.6
GSPO [72] 62.5 72.3 50.8 59.9 66.6 43.5 38.4 34.8 17.0 70.7 72.6 80.6 63.4 71.7
GRPO [45] 61.4 72.7 51.2 59.8 66.8 43.4 39.2 35.9 17.1 70.1 71.1 79.4 61.4 71.0
@@ 4.4 Comparison with State-of-the-arts
We compare our TempR1 against both expert VLP-based methods and recent MLLM-based methods across five temporal understanding tasks, as reported in Tables 1 and 2. As shown in Table 1, TempR1 achieves state-of-the-art performance on Charades-STA [12] and ActivityNet-Caption [22] for TG, as well as on QVHighlights [23] for VHD. Through multi-task joint training, TempR1 attains 61.4 mIoU on Charades-STA, surpassing VideoChat-R1 [27] and MUSEG [37] by 0.6 and 1.7 points, respectively. On the more challenging QVHighlights dataset, which involves multi-segment retrieval, TempR1 shows a significantly larger gain, reaching 71.1 mIoU and outperforming the second-best TAR-TVG [13] by 5.2 mIoU. Remarkably, this performance is achieved after training for only one epoch on the combined multi-task dataset. With additional dataset-specific fine-tuning, TempR1 further improves to 62.5 mIoU on Charades-STA (surpassing TAR-TVG by 1.4 points), 51.9 mIoU on ActivityNet, and 71.6 mIoU on QVHighlights.
As shown in Table 2, TempR1 also excels on more complex tasks, including DTG, GVQA, and TAL. It achieves 59.8 mIoU on ActivityNet [22] for DTG, comparable to the best VLP expert HSCNet [50]. On NExTGQA [62], TempR1 obtains 70.1% question-answering accuracy and 39.2 mIoU for visual evidence localization, improving temporal localization by 3.1 points over VideoChat-R1 [27] while maintaining similar answer accuracy. For TAL on ActivityNet-v1.3 [6], TempR1 achieves 71.0 mF1, outperforming MUSEG [37] by a substantial 13.0 points, validating the effectiveness of our prediction-groundtruth matching strategy and reward design.
We further evaluate TempR1 on general video understanding benchmarks, including VideoMME [11], Perception Test [41], TempCompass [35], and MVBench [51], as illustrated in Figure 3. Using the same training data, SFT tends to weaken general reasoning ability compared with the base model, while our reinforcement fine-tuning notably enhances it. This trend aligns with prior observations in VideoChat-R1 [27] and Time-R1 [56]. Overall, TempR1 delivers the best comprehensive performance across diverse temporal understanding tasks without compromising general video comprehension. These results demonstrate that reinforcement fine-tuning with scalable, task-aligned rewards substantially enhances both temporal comprehension and generalization capabilities of MLLMs in diverse video understanding scenarios.